# The Bayesian Separation Principle for Data-driven Control

### v1

@@ V-A Inference using Gaussian regression
If data are generated according to the linear model (22) and the linear kernel (23) is adopted, one also has
@@
$p(Y)=\frac{\exp\left(-\frac{1}{2}Y^{\top}\Sigma^{-1}Y\right)}{\sqrt{\det(2\pi\Sigma)}}.$ (30)
The above proposition thus provides the predicted output and also its posterior variance through (26) and (27). However, these formulas depend on parameters that may be unknown, particularly the noise variance $\sigma^{2}$ and the kernel scaling factor $\lambda$ (and possibly others like the kernel width in the Gaussian kernel). These variables, often called hyperparameters, can be estimated via the empirical Bayes approach by maximizing the marginal likelihood given by (30) [24]. This tuning approach enjoys important properties. In particular, while the numerator in (30) accounts for data fit, the denominator represents an Occam’s factor which automatically penalizes too complex models [31]. To analyze the structure of the ML estimates it is useful to introduce the concept of degrees of freedom $q(\gamma)$. Using $H(\gamma)$ to denote the so called hat matrix, defined by
@@
As $\gamma$ varies from +$\infty$+ ∞ (strongest regularization) to 0 (absence of regularization), $q(\gamma)$ varies from 0 to the number of data $N$, hence representing a normalized measure of model complexity. After introducing the vector
@@
$\text{WSSU}(\gamma)=\hat{f}^{\top}\mathbb{K}^{-1}\hat{f}$
@@
Combining these two equations, it is then easy to see that the problem can be reduced to finding the zeros of a function with respect to a single unknown variable.
@@
The equality reported in (34) deeply recalls noise variance estimation in linear parametric regression [10] where the degrees of freedom correspond to the dimension of the unknown parameter vector. The correction given by subtracting the number of parameters from the number of data then ensures unbiasedness of the variance estimator.
V-B Asymptotic properties and fast hyperparameters estimation
It is possible to associate each covariance $K$ with a deterministic space of functions known as reproducing kernel Hilbert space (RKHS) denoted by $\mathcal{H}$ [14, 2]. It is generated by all finite and some countably infinite sums of kernel sections [^2] Equivalently, with reference to the expansion (12), the RKHS contains all the sums of functions $\phi_{k}$ with coefficients $\theta_{k}$ satisfying $\sum_{k}\frac{\theta_{k}^{2}}{\zeta_{k}}<\infty$ where, in the stochastic setting, $\zeta_{k}$ represented the variance of $\theta_{k}$.. Any function $g$ in $\mathcal{H}$ can be so written as
$g(x)=\sum_{i}a_{i}K(x_{i},x)\quad\forall x.$
It then follows from (26) that the minimum variance estimate of the system always lies in this space. Instead, before seeing the data, if $K$ includes an infinite number of basis functions any realization of the random field $f$ does not belong to $\mathcal{H}$ almost surely [47]. But even if it has “a priori null probability”, this space is anything but small. The RKHSs induced e.g. by spline and Gaussian kernels enjoy the universality property: they can approximate arbitrarily well any continuous function on every compact [42].
Also in view of this richness, to study the learning rate of a kernel-based estimator, it is customary to assume that the “true” system which has generated the data belongs to $\mathcal{H}$. The asymptotic behaviour of estimates (26) is then analyzed as the data set size $N$ grows to infinity. To some extent, this is a robustness study linked to the model free concept introduced earlier. It investigates if convergence to the true system holds even starting from an infinite-dimensional hypothesis space.
This problem has been studied extensively in statistics and machine learning, see e.g. [7, 8, 45, 48]. However, with some exceptions like [3, 21, 26], results mainly rely on assumptions of independence of the input locations that do not apply in the context of dynamic data. In fact, the $z^{t}$ in (25) contain correlated outputs. The next proposition can be derived from [22] and is particularly suitable for our context. Under reasonable assumptions on the output correlation, it states that (26) converges uniformly to $f$ over any region $X$ that the probability distribution of $z^{t}$ enables to visit. See Appendix for some explanations and additional details on its proof.
Proposition 2
Let the data generated according to
$y(t)=f(z^{t})+e(t)$
where $e$ is any white noise (even non-Gaussian). Consider the spline or Gaussian kernels [^3] The proposition also holds using any polynomial kernel as described in Appendix. The only difference is that it is not possible to state that convergence holds over a set dense in the space of continuous functions since all the RKHSs induced by (21) are finite-dimensional. described in Section IV-C. Let the input locations $z^{t}$ be drawn over the compact $X$ from a probability density function $p$ which assigns non null probability to any open set. Furthermore, for any real-valued $g$, define
[…]
and assume that for any $g$ satisfying
$\|g\|_{p}<q<\infty$
there exists $A$, independent of $g$, such that
$\sum_{k=0}^{\infty}\left|\text{Cov}(g(z^{i}),g(z^{i+k}))\right|<A<\infty$
where Cov is the covariance operator. Then, for a set of systems $f$ in $\mathcal{H}$ dense in the space of continuous functions, if the regularization parameter satisfies
$\gamma\propto N^{1-\alpha},\quad 0<\alpha<\frac{1}{2},$ (35)
as $N$ goes to infinity one has
$\sup_{x\in X}|\hat{f}(x)-f(x)|\longrightarrow_{p}0,$ (36)
where $\longrightarrow_{p}$ denotes convergence in probability.
This result can be seen as the counterpart of the parametric PEM convergence (described at the beginning of the paper) in the nonparametric scenario. We also note that the rule (35) can be combined with (34), hence obtaining a way of tuning the regularization parameter $\gamma$ and the noise variance $\sigma^{2}$ that does not require any iterative method. This point is significant also for MPC cost optimization as illustrated later on.
V-C Explicit expressions of MPC cost
We derive here some explicit expressions of the MPC cost (9), by considering the Gaussian regression framework illustrated before. To simplify notation, we still use $y(t)$ to denote system outputs which are now to be understood as those observed during the control phase of the system (whereas previously they defined elements of $\mathcal{D}$ used during the identification procedure).
First, we observe that the cost (10) becomes
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}\int\|y^{\text{ref}}_{t+\ell}-\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]\|^{2}_{Q_{\ell}}p(Y^{P}_{\ell})dY^{P}_{\ell}$
[…]
$+\sum_{\ell=0}^{T-1}\!\|u^{\text{ref}}_{t+\ell}\!-\!\hat{u}_{t+\ell}\|^{2}_{R_{\ell}}\!$ (37)
where $z_{\ell}^{*}:=[Y^{P}_{\ell}\ y(t-1)\ y(t-2)\ \ldots\ u^{t+\ell}]^{\top}$. We note that the terms $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$ and $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$ in (37) can be computed using Proposition 1, with the dataset $\mathcal{D}$ augmented by $Y^{P}_{\ell}$. In this case, there are Kalman filter-like formulas for updating $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$ and $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$, which have cubic complexity in $T$ (see, e.g., [34]).
In practice, however, since the size of $\mathcal{D}$ is typically much larger than the control horizon $T$, $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$ and $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$ can be approximated by $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$ and $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$, respectively, allowing for the use of the explicit expressions (26) and (27).
The primary challenge in obtaining a closed-form expression of the MPC cost (37) lies in the presence of integrals with respect to $Y_{\ell}^{P}$. In what follows, we explain how this problem can be handled in some relevant situations.
NFIR. When the predictor depends only on past inputs (e.g. the predictor is a FIR system, possibly nonlinear) (37) leads immediately to a closed-form expression of the cost. Indeed, in this case the integrals in (37) w.r.t. $Y_{\ell}^{P}$ disappear and $z_{\ell}^{*}=(u^{t+\ell})^{\top}$.
NARX. When the predictor depends also on past outputs, we can consider the approximation
$\hat{y}^{P}(t)$ $\!=\!f(y(t-1),\ldots,u^{t})$
$\hat{y}^{P}(t+1)$ $\!=\!f(\mathbb{E}[\hat{y}^{P}(t)],y(t-1),\ldots,u^{t+1})$
$\hat{y}^{P}(t+2)$ $\!=\!f(\mathbb{E}[\hat{y}^{P}(t+1)],\mathbb{E}[\hat{y}^{P}(t)],y(t-1),\ldots,u^{t+2})$
[…]
Namely, we replace the conditional future outputs $Y^{P}_{\ell}=[y^{P}(t),\dots,y^{P}(t+\ell-1)]^{\top}$ needed to compute the predictor $\hat{y}^{P}(t+\ell)$ with their estimates $\hat{Y}^{P}_{\ell}=[\mathbb{E}[\hat{y}^{P}(t)],\dots,\mathbb{E}[\hat{y}^{P}(t+\ell-1)]]^{\top}$. With this approximation, the integrals in (37) w.r.t. $Y_{\ell}^{P}$ disappear and $z_{\ell}^{*}=[\hat{Y}^{P}_{\ell}\ y(t-1)\ y(t-2)\ \ldots\ u^{t+\ell}]^{\top}$.

### v2

@@ 5.1 Inference using Gaussian regression
If data are generated according to the linear model (21) and the linear kernel (22) is adopted, the posterior mean and variance of the one-step ahead predictor coefficients in $\theta$ are given by
@@
$p(Y)=\frac{\exp\left(-\frac{1}{2}Y^{\top}\Sigma_{y}^{-1}Y\right)}{\sqrt{\det(2\pi\Sigma_{y})}},\quad\Sigma_{y}=\lambda\Sigma.$ (29)
The above proposition thus provides the predicted output and also its posterior variance through (25) and (26). However, these formulas depend on parameters that may be unknown, particularly the noise variance $\sigma^{2}$ and the kernel scaling factor $\lambda$ (and possibly others like the kernel width in the Gaussian kernel). These variables, often called hyperparameters, can be estimated via the empirical Bayes approach by maximizing the marginal likelihood given by (29) [18]. This tuning approach enjoys important properties. In particular, while the numerator in (29) accounts for data fit, the denominator represents an Occam’s factor which automatically penalizes too complex models [25]. To analyze the structure of the ML estimates it is useful to introduce the concept of degrees of freedom $q(\gamma)$. Using $H(\gamma)$ to denote the so-called hat matrix, defined by
@@
As $\gamma$ varies from $\infty$ (strongest regularization) to 0 (absence of regularization), $q(\gamma)$ varies from 0 to the number of data $N$, hence representing a normalized measure of model complexity. After introducing the vector
@@
$\text{WSSU}(\gamma)=\hat{f}^{\top}\mathbb{K}^{-1}\hat{f}$.
@@
Combining these two equations, the estimate of $\gamma$ solves
$\gamma\frac{\text{WSSU}(\gamma)}{q(\gamma)}=\frac{\text{WSRR}(\gamma)}{N-q(\gamma)}$ (34)
while, for known $\sigma^{2}$, it solves $\gamma\text{WSSU}(\gamma)=\sigma^{2}q(\gamma)$. It then follows that
in Gaussian regression the posterior of $f$, and hence the MPC cost induced by the Bayesian invariance principle, depends only on the regularization parameter $\gamma$ which can be determined simply by solving (34).
Remark 4
@@
The sole dependence on $\gamma$ is due to the fact that, even in the case of unknown noise variance, the latter is linked to $\gamma$ via (33) (and then $\lambda=\sigma^{2}/\gamma$). Note also that (33) deeply recalls noise variance estimation in linear parametric regression [7] where the degrees of freedom correspond to the dimension of the unknown parameter vector. In that setting, the correction given by subtracting the number of parameters from the number of data ensures unbiasedness of the variance estimator.
5.2 Explicit expressions of MPC cost
We derive here some explicit expressions of the MPC cost (9), by considering the Gaussian regression framework illustrated before. We treat separately the case of linear and nonlinear models and, to simplify notation, we let $y^{\text{ref}}_{\tau+\ell}:=y_{\text{ref}}(\tau+\ell)$, $u^{\text{ref}}_{\tau+\ell}:=u_{\text{ref}}(\tau+\ell)$, $\hat{y}^{P}_{\tau+\ell}:=\hat{y}^{P}(\tau+\ell)$, $y_{\tau+\ell}:=y(\tau+\ell)$ $u_{\tau+\ell}:=u(\tau+\ell)$.
5.2.1 Nonlinear case
We begin by analyzing the NFIR case, where the cost (9) can be easily derived in closed form. Then, we consider the NARX case, which requires a more careful analysis due to the dependence on past outputs of the predictors.
NFIR. The output data are generated as
$y(t)=f(u^{t};\theta)+e(t),\\$ (35)
and the conditional predictors $\hat{y}^{P}_{\tau+\ell}=f(u^{\tau+\ell};\theta^{P})$, $\ell=0,\dots,T-1$, depend only on past inputs and $\theta^{P}$.
Since $\hat{y}^{P}_{t+\ell}$ does not depend on $E$, the cost (9) can be written as
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}\!\int\!\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P})p(E)d\theta^{P}dE\!+\!\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$
$=\sum_{\ell=0}^{T-1}\!\int\!\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P})d\theta^{P}\!+\!\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$
$=\sum_{\ell=0}^{T-1}\!\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}]\|^{2}_{Q_{\ell}}\!+\!\sum_{\ell=0}^{T-1}\mathrm{tr}\,Q_{\ell}\ \!\mathrm{Var}[\hat{y}^{P}_{t+\ell}]+\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$ (36)
where in the last step the expectation and variance are taken w.r.t. $\theta^{P}$ and we used the bias-variance decomposition
$\int\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P})d\theta^{P}$
$=\int\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}]-(\hat{y}^{P}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}])\|^{2}_{Q_{\ell}}p(\theta^{P})d\theta^{P}$
$=\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}]\|^{2}_{Q_{\ell}}+\!\underbrace{\int\|\hat{y}^{P}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}]\|^{2}_{Q_{\ell}}p(\theta^{P})d\theta^{P}}_{=\mathrm{tr}\,Q_{\ell}\mathrm{Var}[\hat{y}^{P}_{t+\ell}]}$
$-2(y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}])^{\top}Q_{\ell}\underbrace{\int(\hat{y}^{P}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}])p(\theta^{P})d\theta^{P}}_{=0}.$
In the scalar Gaussian regression framework, we have $\mathrm{tr}\,Q_{\ell}=Q_{\ell}$ and $\mathbb{E}[\hat{y}^{P}_{\tau+\ell}]$, $\mathrm{Var}[\hat{y}^{P}_{t+\ell}]$ are given, respectively, by $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$, $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$ in Proposition 1 with $z_{\ell}^{*}:=[u^{\tau+\ell}]^{\top}$. Thus, the cost (36) admits the closed-form expression
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}Q_{\ell}\,\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]\|^{2}+\sum_{\ell=0}^{T-1}R_{\ell}\,\|u^{\text{ref}}_{\tau+\ell}\!-\!u_{\tau+\ell}\|^{2}$
$+\sum_{\ell=0}^{T-1}\!Q_{\ell}\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]\,.$ (37)
Note that the previous expression is the sum of two terms:
$J_{1}:=\sum_{\ell=0}^{T-1}\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]\|^{2}_{Q_{\ell}}+\sum_{\ell=0}^{T-1}\!\|u^{\text{ref}}_{\tau+\ell}\!-\!u_{\tau+\ell}\|^{2}_{R_{\ell}}$
which can be interpreted as a nominal or “certainty equivalence” cost, and
$J_{2}:=\sum_{\ell=0}^{T-1}\!Q_{\ell}\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$
which acts as a regularizer taking into account the effect on the cost of the uncertainty in predicting future outputs.
NARX. The measurement model is as in (6), where we recall that $y^{t-1}$ and $u^{t}$ are understood as truncated to their first $m$ components.
In this case, the conditional predictors $\hat{y}^{P}_{\tau+\ell}$ depend also on past outputs, and therefore on $E$. It is convenient to replace the dependence of the cost (9) on $\theta^{P}$ and $E$ by a dependence on $\theta^{P}$ and the future conditional outputs $y^{P}_{\tau},y^{P}_{\tau+1},\dots$ contained in $Y^{P}$. This is possible since $(\theta^{P},E)$ and $(\theta^{P},Y^{P})$ generate the same $\sigma$-algebra. As a consequence, letting $Y^{P}_{\ell}:=[y^{P}_{\tau},\dots,y^{P}_{\tau+\ell-1}]^{\top}$, the cost (9) can be written as
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}\!\int\!\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P},Y^{P}_{\ell})d\theta^{P}dY^{P}_{\ell}\!+\!\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$
$=\sum_{\ell=0}^{T-1}\!\int\!\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P}|Y^{P}_{\ell})p(Y^{P}_{\ell})d\theta^{P}dY^{P}_{\ell}\!+\!\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$
$=\sum_{\ell=0}^{T-1}\!\int\!\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[\hat{y}^{P}_{\tau+\ell}|Y^{P}_{\ell}]\|^{2}_{Q_{\ell}}p(Y^{P}_{\ell})dY^{P}_{\ell}+\|u^{\text{ref}}_{\tau+\ell}-u_{\tau+\ell}\|^{2}_{R_{\ell}}$
$+\sum_{\ell=0}^{T-1}\mathrm{tr}\,Q_{\ell}\!\!\int\!\!\mathrm{Var}[\hat{y}^{P}_{t+\ell}|Y^{P}_{\ell}]p(Y^{P}_{\ell})dY^{P}_{\ell}$ (38)
where in the last step the expectation and variance are taken w.r.t. $\theta^{P}$ conditional on $Y^{P}_{\ell}$ and we performed a bias-variance decomposition of the term $\int\!\|y^{\text{ref}}_{\tau+\ell}-\hat{y}^{P}_{\tau+\ell}\|^{2}_{Q_{\ell}}p(\theta^{P}|Y^{P}_{\ell})d\theta^{P}$, similarly as before.
In the scalar Gaussian regression framework, (38) becomes
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}\int\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]\|^{2}_{Q_{\ell}}p(Y^{P}_{\ell})dY^{P}_{\ell}$
$+\sum_{\ell=0}^{T-1}\!Q_{\ell}\!\!\int\!\!\mathrm{Var}[f(z_{\ell}^{*})|\mathcal{D},Y^{P}_{\ell}]p(Y^{P}_{\ell})dY^{P}_{\ell}+\sum_{\ell=0}^{T-1}\!\|u^{\text{ref}}_{\tau+\ell}\!-\!u_{\tau+\ell}\|^{2}_{R_{\ell}}\!$ (39)
where the terms $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$, $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$, with $z_{\ell}^{*}:=[Y^{P}_{\ell}\ y^{\tau-1}\ u^{\tau+\ell}]^{\top}$, can be computed by updating the expressions $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$ and $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$ in Proposition 1 through Kalman filter-like formulas. In fact, to account also for the additional conditioning on $Y^{P}$, one does not have to perform calculations from scratch. The current posterior covariance can be extended over the new measured input locations and the posterior of $f$ can then be updated with complexity cubic in the control horizon $T$, see, e.g., [31]. In practice, however, the size of $\mathcal{D}$ is typically much larger than the control horizon $T$, and $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$, $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D},Y^{P}_{\ell}]$ can be approximated by $\mathbb{E}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$, $\mathrm{Var}[f(z_{\ell}^{*})\,|\,\mathcal{D}]$, respectively, allowing for the direct use of the expressions in Proposition 1.
Beyond this harmless approximation, we introduce the following approximate input locations to obtain a closed-form expression for the cost
$\tilde{z}_{0}^{*}$ $=z_{0}^{*}$
$\tilde{z}_{1}^{*}$ $=[\mathbb{E}[f(\tilde{z}_{0}^{*})\,|\,\mathcal{D}]\ y^{\tau-1}\ u^{\tau+1}]^{\top}$
$\tilde{z}_{2}^{*}$ $=[\mathbb{E}[f(\tilde{z}_{1}^{*})\,|\,\mathcal{D}]\ \mathbb{E}[f(\tilde{z}_{0}^{*})\,|\,\mathcal{D}]\ y^{\tau-1}\ u^{\tau+2}]^{\top}$
$\ \,\vdots$
$\tilde{z}_{T-1}^{*}$ $=[\mathbb{E}[f(\tilde{z}_{T-2}^{*})\,|\,\mathcal{D}]\ \ldots\mathbb{E}[f(\tilde{z}_{0}^{*})\,|\,\mathcal{D}]\ y^{\tau-1}\ u^{\tau+T-1}]^{\top},$
where we stress that the vector
$\mathbb{E}[f(\tilde{z}_{\ell}^{*})\,|\,\mathcal{D}]\ \ldots\ \mathbb{E}[f(\tilde{z}_{0}^{*})\,|\,\mathcal{D}]\ y^{\tau-1}]^{\top},\ \ \ \ell=0,\dots,T-1,$
should be interpreted as truncated at the first $m$ components when used as argument of $f(\cdot)$.
Using the approximate input locations above, the integrals in (39) w.r.t. $Y_{\ell}^{P}$ vanish, and the resulting (approximate) MPC cost is given by
$\hat{J}_{u}$ $=\sum_{\ell=0}^{T-1}Q_{\ell}\,\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[f(\tilde{z}_{\ell}^{*})\,|\,\mathcal{D}]\|^{2}+\sum_{\ell=0}^{T-1}\!R_{\ell}\,\|u^{\text{ref}}_{\tau+\ell}\!-\!u_{\tau+\ell}\|^{2}$
$+\sum_{\ell=0}^{T-1}Q_{\ell}\mathrm{Var}[f(\tilde{z}_{\ell}^{*})|\mathcal{D}]\,.$ (40)
Similarly to the NFIR case, the previous expression can be seen as the sum of a nominal or “certainty equivalence” term
$J_{1}:=\sum_{\ell=0}^{T-1}Q_{\ell}\,\|y^{\text{ref}}_{\tau+\ell}-\mathbb{E}[f(\tilde{z}_{\ell}^{*})\,|\,\mathcal{D}]\|^{2}+\sum_{\ell=0}^{T-1}\!R_{\ell}\,\|u^{\text{ref}}_{\tau+\ell}\!-\!u_{\tau+\ell}\|^{2}$ (41)
and a regularization term
$J_{2}:=\sum_{\ell=0}^{T-1}Q_{\ell}\mathrm{Var}[f(\tilde{z}_{\ell}^{*})|\mathcal{D}]$ (42)
accounting for the uncertainty around the predictors.
5.2.2 Linear case
Consider now systems that are linear in the past inputs and outputs as in (21). For such systems, explicit expressions for both the cost and the optimal input sequence can be derived.